Show the complete solution An anchor chain is wound around a drum with radius 1.5 ft.If the Drum is rotated through an angle of 250° how far will the anchor be lowered?


1
Expert's answer
2021-11-18T17:33:11-0500

The length to which the anchor be lowered is equal to the arc length of rotation on drum.

Arc length"=\dfrac{\theta}{360\degree}\times 2\pi r"

"=\dfrac{250\degree}{360\degree}\times 2\times\dfrac{22}7\times 1.5
\\\approx6.55\ ft"